### v2

# Best practices in Quantum Monte Carlo for metal catalysis: CO hydrolysis on Pt(111)

Abstract
Many scientists are now researching hydrogen production since it affords a clean, sustainable alternative to fossil fuel [4]. Research work and prototyping various aspects of hydrogen power continues to advance.
Over 90 % of all chemical manufacture uses a solid catalyst. This work describes catalytic selective hydrogen production (via water-gas shift, hydration of pre-adsorbed CO) optimising reactant structure on a model catalyst. Focus is on partial O-H bond dissociation after carbon-monoxide is co-adsorbed with water at Pt(111). Finally, hydrogen gas is given off. Many chemical reactions involve bond-dissociation. This process is often the key to rate-limiting reaction steps at solid surfaces. Bond-breaking is poorly described by Hartree-Fock and DFT methods, our embedded active site approach is used. We showcase Quantum Monte Carlo (QMC) methodology using the ground-state Slater Determinant of a simple four primitive-cell layer model, oriented to expose Pt (111), to initialise the QMC. This stochastic approach solves the Schrödinger equation. It recently came of age for heterogeneous systems involving solids. During hydrolysis of carbon monoxide, initial O-H bond stretch is rate-limiting. Its dissociation energy is offset by surface Pt-H bond formation. The reactive formate (H-O-C=O) species formed by initial hydrolysis of CO, also interacting with a vicinal Pt. The products are carbon dioxide and hydrogen, with a H-atom dissociated from the formate species and another desorbed from Pt(111). This yields 100% hydrogen gas.
This single-determinant work is compared to a high-level configuration interaction (CI) wave-function. Activation barriers are given to 2.2kJ/mol (c.f. 0.7 of the CI benchmark). Active sites are embedded in the metal lattice exposing its close-packed face. These trial wave-functions start the QMC.
keywords:
Quantum Monte Carlo calculation, heterogeneous catalysis, metal thin-film surface, low activation barrier
journal: Advances in Quantum Chemistry
1 Introduction
Hydrogen powered trains and cars are now in use and described in the popular press [1]. Jet engines tested on hydrogen by Rolls Royce funded by airlines show the only remaining difficulty is storage requiring 4 times the space of kerosene, and condensation [2]. In the current context of attempting to limit climate change, fossil fuel combustion is scrutinized due to greenhouse gas concerns: hydrogen production has come to the forefront.
Hydrogen can be obtained by electrolysis of water, which requires two O-H bond breakages. The water-gas shift (wgs) reaction is as follows [3]:
$CO+H_{2}O\rightarrow CO_{2}+H_{2}$ (1)
Carbon monoxide (toxic effluent), when hydrated gives hydrogen (N.B. the CO2 can be mineralised as carbonate and removed). CO, pre-adsorbed on metal surfaces, is approached by a physisorbed water molecule. Few metals with modest interaction strength with CO are suitable (e.g. copper interacts too strongly, becoming inert (poisoned), but nickel and platinum are used). Close-packed (111) faces are the most active, when available, because most of these metals have an fcc unit mesh. Adsorbed CO is polar (Cδ+) and reactive towards hydration (unlike free CO). The reaction with water begins by stretching one O-H bond, then the catalyst surface stabilises the product, binding it at a different metal atom. A simple overview is: O=C-(OH)-Pt-Pt-H, where a vicinal Pt atom forms a Pt-H bond. It effectively has much lower activation energy (than water O-H dissociation). Initial water attack of CO at Pt(111) is the rate-limiting step, studied in what follows. Of course, it is not perfectly ’green’ since the catalysts are often based on heavy metals that must be mined and purified. Nevertheless, a catalyst can function without poisoning for an enormous number of cycles (infinite re-use limit). Some can be cleaned and the hydrogen synthesis is a ’grey’ but non-polluting scheme.
Here, platinum catalysts are re-visited. We previously studied them for the water-gas shift reaction [3]. The Pt(111) is the most reactive of cleavages, others are relatively inert. It adsorbs carbon-monoxide, without being poisoned by it. The key hydration step was bench-marked using a sophisticated full configuration-interaction (FCI) wave-function for the embedded active-site reaction intermediate complex. Post Hartree-Fock FCI is time-consuming and scales explosively with system size. Therefore, we now correlate ground-state wave-functions directly, using a generic Jastrow-factor, back-flow and a very careful variational optimisation of the fully converged parameter set. This full parameter set is used for both Transition State (TS) and asymptotic structures involving the same atoms. This strategy has the obvious advantage of ensuring the Slater-Jastrow wave-function will be size consistent. The role of CI is taken over by a generic optimised Jastrow factor. This must be expanded to high order, to converge until coefficients are below a tiny threshold (10-14). It is necessary to use the iso-densic approximation, at least insofar as the compared structures for asymptote and TS comprise the same atoms. Failure to do this can lead to minima that are not size-consistent. The major sources of error, finite-size effect and fixed-node error can be made small. The former is specifically reduced when focussing on surface effects, as studied by Rubenstein et al [37]. This work builds on proof that fixed-node error cancels when resolving energy differences between systems of the same atoms, with few rearranged (i.e. TS and asymptote) [4].
2 Theoretical Method:
Chemical reactions modify electronic correlation energy, at metal catalyst surfaces. Our simple model truncates fcc metals, with a close-packed (111) planar interface. The model compares physisorbed reactants to a transition-state (TS) geometry, located using a Quantum Monte Carlo (QMC) force constant calculation (below). The Transition-State geometry can be optimized using these force constants, as programmed during our earlier work developing the CASINO software [25]. Two structures are defined, involving the same atoms with some bonds modified (water approaching adsorbed CO, on the one-hand, and the TS on the other, for activation barrier calculation). The TS is compact, however, when the water is 7 Å  from the surface (physisorbed), there is a small residual interaction (generally missing from DFT, but accessible to QMC). The TS involves chemisorption that stabilises its product-like form (as a precursor to actual overlap-driven bond formation). A reference active site is defined with a high-level wave-function to start QMC. It caters for reacting molecules and surface atoms describe the metal. This is embedded using periodic potentials. Green’s function methods are applied to heterogeneity of the electronic potential due to the active site [38]. It is input to begin a Quantum Monte Carlo calculation, as follows:
Quantum Monte Carlo proceeds in two stages. First, Variational Monte Carlo (VMC) then, the Diffusion step, Diffusion Monte Carlo (DMC). The final result is almost fully correlated. DMC (and also VMC) scale well with system size (at worst cubically and, in some circumstances linear scaling can be obtained), to efficiently use modern computer architecture that is massively parallel and, above all, produce quantitative, accurate physical properties that are difficult to obtain otherwise. QMC benchmarks of activation barriers for reactions adsorbed on solid catalyst surfaces have such pre-requisites. Some comparisons of gas-phase reactions to their adsorbed counterpart demonstrate a lowering of the barrier, for a given mechanism.
This is already promising, since Density Functional Theory (DFT) often used in these complex systems has been shown to give results opposed to observation. In one example, the reaction appears to be more difficult on the catalyst surface [22]. In other DFT work, as the pseudo-potential improves or the number of electrons approaches the all-electron system, the barrier, initially far below that measured, continues to decrease [23]. These tendencies for DFT results make such an approach unsatisfactory for the bench-marking of barriers. DFT can be used to explore surrounding potential energy surfaces reliably with ab initio or experimental input [24]. Developing and applying QMC calculations on this scale require access to a supercomputer. Since the catalysts are often heavy metals (in this case platinum, Z=78), we need to use pseudo-potentials. Inclusion of semi-core electrons in the valence is certainly necessary when bonding interactions, in the catalytic process involve the transition metal atoms. It also improves test results for platinum [22]. The copper 3d10 shell is dense in the core-region but platinum has its 5d9 6s1 ground-state shell on average further from the core. Pt is thus less prone ’non-locality’ difficulties in pseudo-potentials [25]. Ni has a 4s2 3d8 ground-state and requires semi-core electrons to reduce the non-locality pitfalls observed for copper[25]. They are limited here by using the Casula algorithm [27]. Comparing asymptote and QMC optimised TS geometries for reaction barriers also involves the same atoms, limiting PP errors. Co-ordination of the TS to the metal surface occurs but the distance between atoms is 40 % more than equilibrium bond–lengths and so the role of semi-core electrons is minor. An argon core for Pt is validated [36].
The QMC approach uses statistical physics over a large population, comprising sets of instantaneous particle positions. They are often called ’walkers’ (c.f. the one-dimensional random-walk).
2.1 Input Slater Determinant: Solid-state k-point grids and twist-averaging.
A DFT Generalised Gradient (PBE) calculation over a modest 2x2 k-point grid in the first Brillouin zone can be unfolded in co-ordinate space as QMC input. The product of a Jastrow factor with a Slater determinant gives the trial wave-function (in Slater-Jastrow form). Random numbers actually serve to initialise the ’walkers’ from the trial wave-function giving initial electron density/population). They start a Variational Monte Carlo (VMC) stage. VMC permits trial wave-function optimisation. Its converged output starts the highly accurate Diffusion Monte Carlo (DMC) runs.
Several offset grids (twists) are used for each DMC run, which are then averaged. The poor input is compensated in by relatively long Diffusion Monte Carlo (DMC) runs for 16 twists.
2.2 Variation Monte Carlo:
A preliminary VMC calculation is carried out in order to generate several thousand configurations (instantaneous positions i.e. electron co-ordinates). VMC is driven by minimisation. The so called local energy is minimised:
$\frac{H\psi}{\psi}$ (2)
The kinetic energy terms involved are smoother and have lower variance in exponentially decaying bases, as shown in our work on wave-function quality [30]. A Jastrow factor including electron pair, electron-nucleus and three-body (two-electron and nucleus) is defined. This Jastrow factor [28, 29] is carefully optimised with respect to its polynomial coefficients (up to 5 % of the total time). This factor uses a polynomial expansion in the variables of explicit correlation.
2.2.1 Generic Jastrow factor
A Slater-Jastrow wave-function is defined as the product of $\Psi_{S}$, a Slater determinant and the exponential with an explicit correlation Jastrow factor J(R) as argument:
$\Psi{\bf(R)}=e^{J(R)}\Psi_{S}{\bf(R)}=e^{J{\bf(R)}}D_{u}{\bf(R_{u})}\,D_{d}{\bf(R_{d})}$ (3)
The determinants $D_{u}(R_{u})$ and $D_{d}(R_{d})$ treat ’up’ and ’down’ spin separately if necessary. This new approach [29] which generates a huge parameter set was tested for our platinum slab model, introduced above and detailed below. The set was converged in its full extent, but can be reduced somewhat by symmetry (grouping atoms). Here, the DMC steps are rather time-consuming.
The generic Jastrow factor used (in Common Algebraic Specification Language (CASL)), had polynomials in e-e, e-n (to order 9) and e-e-n (to order 4), with cut off radius 10 au.
Truncation to a finite range involves a factor:
$t(R)=(r-L)^{C}\Theta_{H}(r-L)$ (4)
Take C=3, since the wave-function and at least two derivatives are continuous. $\Theta_{H}$ is the Heaviside step-function (Wigner Seitz cell diameter L).
$J({\bf R})=\sum_{i<j}^{N}u_{P_{ij}}(r_{ij})\,+\,\sum_{i}^{N}\,\sum_{I}^{N_{n}}\chi_{S_{iI}}(r_{iI})\,+\,\sum_{i<j}^{N}\sum_{I}^{N_{n}}f_{T_{ijI}}(r_{iI},r_{jI},r_{ij})$ (5)
$u,\chi$ and $f$ are polynomials (with variational coefficients) in the inter-particle distances. The symbols P, S, T denote the channel indices. The $u$ term caters for inter-electron contributions over $r_{ij}$ and the $\chi$ term for electron-nuclear terms over $r_{iI}$. They both obey Kato’s cusp conditions. The $f$ term is the three-particle electron pair and nucleus contribution (over $r_{ij}$ and two $r_{iI}$ values). This 3-body $f$ term cannot be neglected. Values of $f$ evolve in Jastrow parameter-set optimisations, and compensate the simpler active site wave-functions used. This 3-body term involves ’double-counting’, some contributions already included in electron-electron and electron-nucleus terms. It is avoided when cutoff functions are switched on [31].
In the present work, DMC runs are short, because the Jastrow factor is fully optimised, and we have obtained convergence for variance followed by energy Variation Monte Carlo (VMC). High-order expansions in the instantaneous inter-particle distance (9 for 2 particle an 4 for 3 particle) lead to large parameter sets (550 here), since individual atoms are treated. When all unbound Pt-atoms are grouped, and also with 3rd order f-expansion, VMC converges faster, but to higher energy. It is therefore an advantage to have converged the complete Jastrow factor, and operate with minimum energies.
Occasional large DMC populations and sub-variational data over portions of DMC runs are due to pseudo-potential non-locality. To remove this defect, non-local grid energy is evaluated. For 3d electrons it is greater than for higher d-shells (Ni is a 3d transition metal, see [25]). Pt is not a 3d element, nevertheless, a specific high-order 7 integration grid is used for non-local terms here. These systems involving stretched bonds need accurate evaluation of electronic correlation (not just differences between energy values). QMC has proven to be a reliable method for describing stretched bonds [33]. Generic Jastrow factors were used for this molecule-solid system (with 21 terms, one per atom). Whilst two surface inter-layer distances are optimised, incorportating the embedded structure, bulk platinum periodicity is resumed beyond the 2nd atom-layer. Note that $z\leq 0$ is perpendicular to the thin film Pt(111) surface. The x and y dimensions should be much longer than the maximum bond-length to limit finite-size effects. The surfaces in this model are planar, whereas it is known that metal close-packed faces re-arrange forming high Miller index ridges or meanders [34]. Our planar-surface model uses the transition state (TS) and a reference state with the same atoms as the TS, before the reaction.
2.3 Diffusion Monte Carlo:
During DMC runs, the ground-state component of the trial wave function is projected out by solving the Schrödinger equation (SWE) in imaginary time. DMC is carried out in the fixed-phase approximation, which uses the nodes from the input trial wave-function. An updated overview of the VMC and DMC methods is given in [19]. These trial wave-functions can be optimised with a complex Jastrow factor, [26] because they potentially provide the input with exact nodes. This improves the single Slater determinant describing a ground-state from the DFT orbitals for heterogeneous systems. DFT ground-state orbitals decay exactly, but Jastrows are needed to improve its nodes. They are also improved by embedding the MRCI active site [10].
The imaginary-time SWE is transformed into a Diffusion equation in the 3N-dimensional space of electron coordinates (with the potential energy acting as a source/sink term) by replacing $t$ by the pure-imaginary time $it$:
The time-dependent Schrödinger equation may be written as:
$i\,\frac{d}{dt}\Psi(r,t)=-\frac{1}{2}\nabla^{2}\Psi({\bf{r}},t)\,+\,V({\bf{r}})\Psi({\bf{r}},t)$ (6)
where $\,\nabla^{2}$ is the scalar Laplacian of an instantaneous electron-position vector: $\,{\bf{r}}$. Defining the Hamiltonian (for stationary states):
$\mathcal{H}=-\frac{1}{2}\nabla^{2}\,+\,V({\bf{r}})$ (7)
[…]
$-\,\frac{d}{d\tau}\Psi({\bf{r}},t)=(\mathcal{H}-E_{0})\,\Psi({\bf{r}},t)$ (8)
The imaginary-time SWE can therefore be solved by combining diffusion and branching or dying processes. Introducing importance sampling, using the trial wave function transforms the problem into one involving drift as well as diffusion but greatly reduces the population fluctuations due to the branching/dying process. The Fermionic wave function antisymmetry is maintained by constraining the nodal surface to equal that of the trial wave function. To resolve energy values, QMC statistical error in total values must be kept small. Initially, a 0.005 au time-step is used, extrapolated to the limit 0 to account for possible time-step bias. Assuming the equilibration of a QMC calculation is of negligible duration, $\sigma$ falls off as 1/$\sqrt{T}$, where T is the cpu-time (h). Accuracy for a given time also depends on the quality (nodes) of the trial wave-function. Lower fixed-phase error reduces equilibration time.
Fixed-node error (fixed phase, of complex wave-functions) is systematic and steps must be taken to make it small. Single-determinant DFT ground-states are used for heteorgeneous systems here. Their influence is limited, however, by evaluating energy differences between structures with the same atoms. The imaginary-time SWE is transformed into a Diffusion equation in the 3N-dimensional space of electron coordinates. Convergence of the DMC properties is quadratic, therefore, energy eigen-values may converge from above or below the exact value. The DMC algorithm can be made variational in the Casula T-move scheme [27], which was used in this work.
3 Models
We return to focus on water-gas shift hydrogen production [21]. Our reference work, for bulk platinum, shows that a supported Pt (111) monolayer is more reactive than bulk platinum, limited by Pt (111). These studies describe water attack on pre-adsorbed CO at Pt(111). We found the activation barrier for this step on bulk Pt 71 ±$\pm 0.7$ kJ/mol in [10, 9] and obtained a reliable estimate for the monolayer supported by Al(111) at 64.5 kJ/mol.
This section requires several comments on the model catalyst. Defining a supercell that is 4 mesh thick compares well to the benchmark (using a 5 mesh thick slab for identical top and bottom), because of periodic continuation in the solid. This procedure is simply more convenient with multiples of 4, which is thick enough to avoid surface (skin) 2D-state effects.
Figure 1: Primitive unit cell for Pt, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  CO is adsorbed in Pt3 hollows.
One author consistently used Quantum Monte Carlo methods at chemical accuracy ( of 1 kcal/mol = 4.2 kJ/mol) for activation barriers of reactions at metal catalyst surfaces, since publishing test benchmarks on hydrogen dissociation on Cu(111) in 2015. Progress has been made in controlling error sources with the Casino QMC code for periodic systems (Doblhoff-Dier and Hoggan [25, 20]). Our recent publication gives the associated rate-limiting barrier to within 1kJ/mol [10, 9]. This benchmark uses a Multi-reference (MRCI) trial wave-function, suited to bond breaking and formation and an embedded active site approach to the catalyst, allowing high-level input to be used for a small molecular active site, before the periodic QMC. Other applications of QMC to catalysis and weak interactions of accuracy close to 1 kJ/mol are rare and often used to calibrate DFT functionals for cheaper investigations [12, 13, 14]. The above reference asymptotic guiding wavefunction is obtained for the structure below, and simply taken from our previous work [10, 9]. Reliable DFT results on catalytic hydrogen production are [17, 18]. In this work, we compare the TS, from single determinant input to the multi-reference benchmark.
Figure 2: Primitive unit cell for Ni, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  TS closeup (note H-atom at x=y=0.
It would be interesting to obtain valid information on the catalytic process with simpler wave-functions, if possible. Using single-determinant (ground state) wave-functions is assessed here, and already tested for CO on Pt(111) [15]. Limiting systematic error due to imperfect nodes from simpler trial wave-functions to 2.5 kJ/mol was studied in [15]. Barriers are over-estimated since transition states are strongly multi-reference. Embedding proceeds smoothly using a Green’s matrix addition of electronic potentials [16, 38].
3.1 The stages of our approach are as follows:
1-Define the active site with 2 Pt-atoms in the (111) face. One receives H the other describes the equilateral triangle motif, using translational symmetry. The hollow site where CO is adsorbed is at this triangle’s centre of mass: The atoms involved are Pt1, Pt2, CO and H2O. TS input geometry, from QMC force constants (see below).
Tabulated active site co-ordinates follow (a.u.): Atom x y z Pt1 0.0000 0.0000 0.0000 Pt2 2.6137 4.5271 0.0000 C 2.6137 3.0179 4.3000 O 2.6137 4.4344 6.5864 O 1.9000 0.0000 4.3000 H 0.0000 0.0000 3.0500 H 1.3155 0.0000 6.0126 2-Molpro [35] uses this input to evaluate an MRCI wave-function as in [10, 9]., which is embedded in the Pt-lattice with a 2x2 atom face and 5 layers; ABABA. This defines a periodic slab. The ’active site’ embedded in Pt-slab (without Pt-atom duplication) is expanded in plane waves, with periodic boundary conditions in k-space. Embedding simply uses standard Green matrix addition for the electronic potential [38]. In this work, comparison to a single determinant input is studied, to assess its effect on accuracy.
2-Resulting Slater determinant wave-functions are then expressed in Slater-Jastrow form (Kohn-Sham determinant and Jastrow function product).
3-This trial wave function contains arbitrary parameters (polynomial coefficients of inter-particle distances in the Jastrow factor) that are optimised using Variational Monte Carlo (VMC).
4-The resulting optimised trial wave function is used to generate initial configurations variationally for the Diffusion Monte Carlo step (DMC). Pseudo-potentials have been discussed in previous work [25] and require a brief recap in the section below.
This wave-function is then used to generate a population of ’walkers’ (configurations) that are propagated in imaginary time during the second, diffusion step (DMC). DMC uses the fixed-node approximation, conserving the nodes on an input trial wave-function. CASINO software [19] is used although it is not ’black box’ software since only operator choices lead to efficient Jastrows.
A periodic solid is truncated by a planar close-packed platinum surface. The toxic CO is used to selectively produce hydrogen in Pt(111) catalysed water-gas shift processes. The active site-reaction complex is embedded into this face. It comprises the optimal TS for pre-adsorbed CO hydrolysis. We approach the carbonyl carbon (adsorbed at the platinum surface, bound to the atom denoted Pt1) with the water oxygen, whilst adsorbing the dissociating H-atom on a vicinal Pt2 atom. The TS is defined as an energy saddle-point, located by minimising all co-ordinates except the reaction co-ordinate which is the C-carbonyl Water O distance, in a weighted average with the O-H stretch in water.
The geometry has been confirmed determining force-constants to identify the reaction-pathway, using algorithms developed in [32] which is now available for QMC. From this geometry, the reactants or products are obtained by steepest descent energy minimisation.
4 Results
The asymptote is taken from our benchmark, since it is the TS that can be expected to show strong multi-reference character, and we test the infuence of this on activation barrier accuracy by comparing it with single-determinant work conducted in the present study. Here, only $4000$ DMC steps for each geometry converge it, to an error in standard error (se) ratio $R<1/10$. Furthermore, the grid-sizes used are modest (2 2 1), compared with converged DFT grids of 16 16 1. Converged DFT total energies are used as control-variate in the QMC twist-averaging process [19], using 4 twists at present and time-step $\tau$ = 0.005 au.
Defining ”corr” as time-correlation, given as a number of steps. E in au. $4200$ steps; corr=262 use the best reblock length (256) standard error (se) = 1.583 mHa/ error in se 0.3066 mHa
Ets = $-462.035595$ approx. -462.0356 $\bf(max-462.0340)$
$2770$ steps; corr gives $32$ 1.03/32 Ets = -462.03324 $\bf(max-462.03424$   similar to above)
Best
$-462.0356\pm 1.6mHa$
Asymptote (referenvce value) [10]
$-462.0629\pm 0.2mHa$
BARRIER = $71.2$ kJ/mol ±$\pm 4.2$ kJ/mol This model reaches the threshold of chemical accuracy, in spite of limited resources required (1 kcal/mol = 4.18 kJ/mol). The error bars on the TS are 5 times those obtained using MRCI and FCI embedded active site input, i.e. compared to the benchmark [10]. Nevertheless, the activation barrier agrees with the 71 ±$\pm 0.7$ kJ/mol obtained in the benchmark. They are therefore consistent, but could be several kJ/mol apart.
A further run, extrapolated to the zero time-step (tau, $\tau\rightarrow 0$) limit converges well, with a correlation length just over 64 steps (reblocked as 64), for which the accuracy of the barrier is $\pm$ 2kJ/mol. This is continued for a total of 6000 steps and the onset of long-range oscillatory behaviour is observed, leading to an amplitude of 1.5 kJ/mol which is within the variance obtained by twist-averaging over 4 twists. $5600$ steps; corr=82 use best block (64) $64$ converged se is 0.93 mHa (error in standard error 0.0735 mHa), less than R = 1/10 Ets = $-462.036720$ (max -462.0360) absolute difference with $\tau$ = 0.005 runs is 3kJ/mol or 2 standard errors, so not beyond Gaussian significance.
taking the reference asymptote: EXTRAPOLATED $\tau$ BARRIER: = $68.8$ kJ/mol ±$\pm 2.2$ kJ/mol Time-step bias is removed by the extrapolation of steps to zero. Accuracy is improved, by a factor two, remaining more than twice that of the benchmark. All thtree barrier estimates are well within 2kJ/mol and the ’extrapolated tau’ mark has a variationally lower-better TS). The barrier agrees with the 71kJ/mol obtained in the benchmark in [10], however, it has twice the error bar. They are consistent (up to a couple of kJ/mol apart). Finally, we can quote the activation barrier, with a significantly narrowed 95 % confidence interval (which also betters the long calculations of [15]) as: Ea = 70.2 kJ/mol ±$\pm 2.2$ kJ/mol This is consistent with all the numerical results of this work, and close to the benchmark at 71 ±$\pm 0.7$ kJ/mol for the rate-limiting barrier of CO hydrolysis on Pt(111) remains our reference.
The 4 mesh thick slab of section 3, co-ordinates adsorbed molecules above ithe (111) face. It has also become clear that these DMC runs include a sytematic error, also related to the 4-mesh Platinum catalyst model. It stems from the Pt-pseudo-potential that has significant non-locality [25] and was contained by high order non-local grid integration. Convergence tests show non-local PP error in total energy is $\leq 1.5$ kJ/mol.
Despite this, DMC populations of walkers fluctuate by at most 5-10 % and the R threshold for converged runs at 1/10 is reached by 5 600 data-points without significant time-step bias and blocks of 256 points in averaging. Fixed-node error is also present, however it can be made small, even with the present Single-determinant wave-functions [26]. This is represents a metal surface and the reaction is conducted realistically. the bulk solid underneath is periodic.
5 Conclusions and perspectives
This work uses a simple slab of four Pt4 unit-mesh describing a (111) surface, adsorbing C=O, with an unperturbed bulk lattice below it. Finite-size errors have been studied and shown to be small for surfaces [37]. There are, nevertheless signs of fixed-node error, that does not entirely cancel by comparing energies of reactants and a TS containing the same atoms.
A semi-quantitative conjecture, which we have previously observed in ’proof of concept’ QMC on metal catalysts (see [15]) is that fixed-node error will have an upper bound of 10% of the total, for such energy differences. This conjecture stems from tests of Casino software by Needs and Towler. It is hence desirable to use CI wave-function input, that will be a subject of future work. The Pt (5d9 6s1) pseudo-potential has some non-locality [25]. No redox phenomena are observed, therefore, this partly occupied 3d-shell is necessarily valence. Non-locality effects have been kept below 2kJ/mol for all geometries studied.
Taking the median barrier, between the zero time-step extrapolation estimate and the runs at a time-step of 0.005 au, we see that all the variances and absolute differences fall in the range:
Ea = 70 $\pm$ 2 kJ/mol close to the benchmark value of 71 $\pm$ 0.7 kJ/mol
This time-saving strategy, using a large parameter set Generic Jastrow, for all atoms, when correlating the guiding wave-function of the 4 mesh slab with adsorbates (21 atoms in all) is therefore very promising. In particular it cuts cpu time by a factor of 14 from the slightly less-accurate [15].
Acknowledgements. We thank GENCI for a special allocation on Irene (nov 2022 to nov 2023 CEA, Bruyères-le-Châtel, F) and renewals until 2025, as well as University of Clermont Mesocentre (data preparation, tests).

### v3

@@
 Abstract
-Many scientists are now researching hydrogen production since it affords a clean, sustainable alternative to fossil fuel [4].
+Many scientists are now researching hydrogen production, because it provides a clean, sustainable alternative to fossil fuel [4].
 Research work and prototyping various aspects of hydrogen power continues to advance.
 Over 90 % of all chemical manufacture uses a solid catalyst.
-This work describes catalytic selective hydrogen production (via water-gas shift, hydration of pre-adsorbed CO) optimising reactant structure on a model catalyst.
+This work describes catalytic selective hydrogen production (hydration of pre-adsorbed CO, during the water-gas shiftreaction) optimising reactant structure on a model catalyst.
 Focus is on partial O-H bond dissociation after carbon-monoxide is co-adsorbed with water at Pt(111).
@@
 Its dissociation energy is offset by surface Pt-H bond formation.
-The reactive formate (H-O-C=O) species formed by initial hydrolysis of CO, also interacting with a vicinal Pt.
+The reactive formate (H-O-C=O) species is formed, and interacts with a vicinal Pt.
 The products are carbon dioxide and hydrogen, with a H-atom dissociated from the formate species and another desorbed from Pt(111).
@@
 The key hydration step was bench-marked using a sophisticated full configuration-interaction (FCI) wave-function for the embedded active-site reaction intermediate complex.
-Post Hartree-Fock FCI is time-consuming and scales explosively with system size.
-Therefore, we now correlate ground-state wave-functions directly, using a generic Jastrow-factor, back-flow and a very careful variational optimisation of the fully converged parameter set.
+Accurate post Hartree-Fock FCI is time-consuming and scales explosively with system size.
+Therefore, we now correlate ground-state wave-functions directly, using a generic Jastrow-factor (gJF), back-flow and a very careful variational optimisation of the fully converged parameter set.
 This full parameter set is used for both Transition State (TS) and asymptotic structures involving the same atoms.
-This strategy has the obvious advantage of ensuring the Slater-Jastrow wave-function will be size consistent.
-The role of CI is taken over by a generic optimised Jastrow factor.
-This must be expanded to high order, to converge until coefficients are below a tiny threshold (10-14).
+The advantage of this strategy lies in ensuring the Slater-Jastrow wave-function will be size consistent.
+The role of CI is taken over by a generic optimised gJF.
+The gJF must be expanded to high order, to converge until coefficients smaller than, say (10-14).
 It is necessary to use the iso-densic approximation, at least insofar as the compared structures for asymptote and TS comprise the same atoms.
 Failure to do this can lead to minima that are not size-consistent.
-The major sources of error, finite-size effect and fixed-node error can be made small.
-The former is specifically reduced when focussing on surface effects, as studied by Rubenstein et al [37].
-This work builds on proof that fixed-node error cancels when resolving energy differences between systems of the same atoms, with few rearranged (i.e.
-TS and asymptote) [4].
+Main sources of (non-statistical) error, finite-size effect and fixed-node error can be made small.
+The former is specifically reduced when focussing on surface effects, as shown by Rubenstein et al [37].
+We use the fact that fixed-node error cancels when evaluating energy differences between systems comprising the same atoms, with few rearranged (TS and asymptote) [4].
 2 Theoretical Method:
 Chemical reactions modify electronic correlation energy, at metal catalyst surfaces.
-Our simple model truncates fcc metals, with a close-packed (111) planar interface.
-The model compares physisorbed reactants to a transition-state (TS) geometry, located using a Quantum Monte Carlo (QMC) force constant calculation (below).
-The Transition-State geometry can be optimized using these force constants, as programmed during our earlier work developing the CASINO software [25].
-Two structures are defined, involving the same atoms with some bonds modified (water approaching adsorbed CO, on the one-hand, and the TS on the other, for activation barrier calculation).
+Fcc metals are truncated with a close-packed (111) planar interface.
+The model compares physisorbed reactants to a transition-state (TS) geometry, located using a Quantum Monte Carlo (QMC) force constant calculation.
+Transition-State nuclear positions can be optimized using these force constants, we programmed for the CASINO software [25].
+QMC is then applied to the electrons, for ’clamped’ nuclei.
+Two structures are defined, involving the same atoms with some bonds modified (water approaching adsorbed CO, and the hydration TS, for activation barrier calculation).
 The TS is compact, however, when the water is 7 Å  from the surface (physisorbed), there is a small residual interaction (generally missing from DFT, but accessible to QMC).
@@
 It also improves test results for platinum [22].
-The copper 3d10 shell is dense in the core-region but platinum has its 5d9 6s1 ground-state shell on average further from the core.
-Pt is thus less prone ’non-locality’ difficulties in pseudo-potentials [25].
-Ni has a 4s2 3d8 ground-state and requires semi-core electrons to reduce the non-locality pitfalls observed for copper[25].
-They are limited here by using the Casula algorithm [27].
+The copper 3d10 shell is dense in the core-region and delocalised in the conduction bands, but platinum has its 5d9 6s1 ground-state shell on average further from the core.
+Cu and Ni (with a 4s2 3d8 ground-state) are prone to ’non-locality’ effects, in pseudo-potentials [25].
+Pt is less affected.
+All require semi-core valence shells [25].
+These effects are limited by the Casula T-move algorithm [27] (in DMC).
 Comparing asymptote and QMC optimised TS geometries for reaction barriers also involves the same atoms, limiting PP errors.
@@
 Here, the DMC steps are rather time-consuming.
-The generic Jastrow factor used (in Common Algebraic Specification Language (CASL)), had polynomials in e-e, e-n (to order 9) and e-e-n (to order 4), with cut off radius 10 au.
+The generic Jastrow factor used (in Common Algebraic Specification Language (CASL)), comprised polynomials in e-e, e-n (to order 9) and e-e-n (to order 4), with cut off radius around 10 a.u.
+(see aapendix).
 Truncation to a finite range involves a factor:
@@
 It is avoided when cutoff functions are switched on [31].
-In the present work, DMC runs are short, because the Jastrow factor is fully optimised, and we have obtained convergence for variance followed by energy Variation Monte Carlo (VMC).
-High-order expansions in the instantaneous inter-particle distance (9 for 2 particle an 4 for 3 particle) lead to large parameter sets (550 here), since individual atoms are treated.
+In the present work, DMC runs require few data-points, the Jastrow factor being fully optimised, and we have obtained convergence for variance followed by energy Variation Monte Carlo (VMC).
+High-order expansions in the instantaneous inter-particle distance (9 for 2 particle an 4 for 3 particle) give large parameter sets (550 here with individual atoms are treated).
 When all unbound Pt-atoms are grouped, and also with 3rd order f-expansion, VMC converges faster, but to higher energy.
@@
 DFT ground-state orbitals decay exactly, but Jastrows are needed to improve its nodes.
-They are also improved by embedding the MRCI active site [10].
+They are also improved by embedding the Multi-reference (MRCI) active site [10].
 The imaginary-time SWE is transformed into a Diffusion equation in the 3N-dimensional space of electron coordinates (with the potential energy acting as a source/sink term) by replacing $t$ by the pure-imaginary time $it$:
@@
 Other applications of QMC to catalysis and weak interactions of accuracy close to 1 kJ/mol are rare and often used to calibrate DFT functionals for cheaper investigations [12, 13, 14].
-The above reference asymptotic guiding wavefunction is obtained for the structure below, and simply taken from our previous work [10, 9].
+The above reference asymptotic guiding wavefunction is obtained for the structure in Figure 1, and simply taken from our previous work [10, 9].
 Reliable DFT results on catalytic hydrogen production are [17, 18].
-In this work, we compare the TS, from single determinant input to the multi-reference benchmark.
-Figure 2: Primitive unit cell for Ni, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  TS closeup (note H-atom at x=y=0.
+In this work, we compare the TS (the strcture is optimised with QMC forces and plotted in Figure 2), from single determinant input to the multi-reference benchmark.
+Figure 2: Primitive unit cell for Ni, exposing (111) surfaces top and bottom. The lattice parameter a, is 3.92Å  TS closeup (note H-atom at x=y=0.)
 It would be interesting to obtain valid information on the catalytic process with simpler wave-functions, if possible.
@@
 In this work, comparison to a single determinant input is studied, to assess its effect on accuracy.
-2-Resulting Slater determinant wave-functions are then expressed in Slater-Jastrow form (Kohn-Sham determinant and Jastrow function product).
-3-This trial wave function contains arbitrary parameters (polynomial coefficients of inter-particle distances in the Jastrow factor) that are optimised using Variational Monte Carlo (VMC).
-4-The resulting optimised trial wave function is used to generate initial configurations variationally for the Diffusion Monte Carlo step (DMC).
+3-Resulting Slater determinant wave-functions are then expressed in Slater-Jastrow form (Kohn-Sham determinant and Jastrow function product).
+4-This trial wave function contains arbitrary parameters (polynomial coefficients of inter-particle distances in the Jastrow factor) that are optimised using Variational Monte Carlo (VMC).
+5-The resulting optimised trial wave function is used to generate initial configurations variationally for the Diffusion Monte Carlo step (DMC).
 Pseudo-potentials have been discussed in previous work [25] and require a brief recap in the section below.
@@
 4 Results
-The asymptote is taken from our benchmark, since it is the TS that can be expected to show strong multi-reference character, and we test the infuence of this on activation barrier accuracy by comparing it with single-determinant work conducted in the present study.
+To test the validity of activation barriers from QMC with single-determinant input, the asymptote is taken from our benchmark, since it is the TS that can be expected to show strong multi-reference character.
+The benchmark activation barriers are compared with single-determinant work conducted in the present study.
 Here, only $4000$ DMC steps for each geometry converge it, to an error in standard error (se) ratio $R<1/10$.
 Furthermore, the grid-sizes used are modest (2 2 1), compared with converged DFT grids of 16 16 1.
-Converged DFT total energies are used as control-variate in the QMC twist-averaging process [19], using 4 twists at present and time-step $\tau$ = 0.005 au.
-Defining ”corr” as time-correlation, given as a number of steps. E in au. $4200$ steps; corr=262 use the best reblock length (256) standard error (se) = 1.583 mHa/ error in se 0.3066 mHa
-Ets = $-462.035595$ approx. -462.0356 $\bf(max-462.0340)$
-$2770$ steps; corr gives $32$ 1.03/32 Ets = -462.03324 $\bf(max-462.03424$   similar to above)
-Best
-$-462.0356\pm 1.6mHa$
-Asymptote (referenvce value) [10]
-$-462.0629\pm 0.2mHa$
-BARRIER = $71.2$ kJ/mol ±$\pm 4.2$ kJ/mol This model reaches the threshold of chemical accuracy, in spite of limited resources required (1 kcal/mol = 4.18 kJ/mol).
-The error bars on the TS are 5 times those obtained using MRCI and FCI embedded active site input, i.e.
-compared to the benchmark [10].
+Converged DFT energy values are the control-variate in our QMC twist-averaging process [19], using 4 twists here, and time-step $\tau$ = 0.005 au.
+Tabulated averaged raw QMC runs (a.u.): Defining ”corr” (time-correlation) as a number of steps.
+The best reblock length, is labeled ’rl’.
+The standard error (se) and E, are in Hartree (Ha).
+Steps corr best rl se mHa error in se Ets a $4200$ 262 256 1.583 0.3066 $-462.035595$ b $2770$ 34 32 1.03 0.32 -462.03324 Tabulated analysis of error-strewn barrier height, in Hartree (Ha) and standard error (se) units specified: Run.
+Etot maximum se mHa Barrier kJ/mol se kJ/mol a $-462.0356$ $-462.0340$ 2 - - b $-462.0332$ $-462.03424$ 2 - - best $-462.0356$ 1.6 - - asymptote $-462.0629$ 0.2 $71.2$ $4.2$ benchmark - - 0.2 $71$ $0.7$
+This model reaches the threshold of chemical accuracy, in spite of limited resources required (1 kcal/mol = 4.18 kJ/mol).
+The TS standard error is 5 times that of the benchmark [10], obtained using MRCI (FCI) embedded active site input.
 Nevertheless, the activation barrier agrees with the 71 ±$\pm 0.7$ kJ/mol obtained in the benchmark.
@@
 This is continued for a total of 6000 steps and the onset of long-range oscillatory behaviour is observed, leading to an amplitude of 1.5 kJ/mol which is within the variance obtained by twist-averaging over 4 twists.
-$5600$ steps; corr=82 use best block (64) $64$ converged se is 0.93 mHa (error in standard error 0.0735 mHa), less than R = 1/10 Ets = $-462.036720$ (max -462.0360) absolute difference with $\tau$ = 0.005 runs is 3kJ/mol or 2 standard errors, so not beyond Gaussian significance.
-taking the reference asymptote: EXTRAPOLATED $\tau$ BARRIER: = $68.8$ kJ/mol ±$\pm 2.2$ kJ/mol Time-step bias is removed by the extrapolation of steps to zero.
+Time-step bias is removed by the extrapolation of steps to zero.
 Accuracy is improved, by a factor two, remaining more than twice that of the benchmark.
@@
 They are consistent (up to a couple of kJ/mol apart).
-Finally, we can quote the activation barrier, with a significantly narrowed 95 % confidence interval (which also betters the long calculations of [15]) as: Ea = 70.2 kJ/mol ±$\pm 2.2$ kJ/mol This is consistent with all the numerical results of this work, and close to the benchmark at 71 ±$\pm 0.7$ kJ/mol for the rate-limiting barrier of CO hydrolysis on Pt(111) remains our reference.
-The 4 mesh thick slab of section 3, co-ordinates adsorbed molecules above ithe (111) face.
-It has also become clear that these DMC runs include a sytematic error, also related to the 4-mesh Platinum catalyst model.
-It stems from the Pt-pseudo-potential that has significant non-locality [25] and was contained by high order non-local grid integration.
+After longer runs and extrapolation to zero time-step, we again take the reference asymptote.
+The absolute difference with $\tau$ = 0.005 runs is 3kJ/mol or 2 standard errors, so not beyond Gaussian significance.
+Tabulated barriers from longer runs (refined barrier heights) Again, ”corr” is time-correlation, given as a number of steps.
+The best reblock length is labeled ’rl’.
+The se (standard error) and E, are in Hartree steps Ets max.
+corr/rl se mHa err.
+in se Ea kJ/mol a $5600$ $-462.03672$462.036 84 / 64 0.93 0.0735 $68.8$$\pm 2.2$
+This is consistent with all the numerical results of this work, and close to the benchmark at 71 ±$\pm 0.7$ kJ/mol for the rate-limiting barrier of CO hydrolysis on Pt(111) remains our reference.
+Finally, we can quote the activation barrier, with a significantly narrowed 95 % confidence interval (which also betters the long calculations of [15]) as: Ea = 70.2 kJ/mol ±$\pm 2.2$ kJ/mol The 4 mesh thick slab (see section 3) has adsorbed molecules above the (111) face.
+These DMC runs include sytematic error, due to the 4-mesh Platinum catalyst model.
+It stems significant non-locality [25] in the Pt pseudopotentil, and was contained by high order non-local grid integration.
 Convergence tests show non-local PP error in total energy is $\leq 1.5$ kJ/mol.
-Despite this, DMC populations of walkers fluctuate by at most 5-10 % and the R threshold for converged runs at 1/10 is reached by 5 600 data-points without significant time-step bias and blocks of 256 points in averaging.
+Having shown this ’non-locality’ PP-defect persists in the DMC runs (gently oscillating energy values), causing the populations of its walkers fluctuate by at most 5-10 % and the R threshold for converged runs at 1/10 is reached by 5 600 data-points without significant time-step bias and blocks of 256 points in averaging.
 Fixed-node error is also present, however it can be made small, even with the present Single-determinant wave-functions [26].
@@
 In particular it cuts cpu time by a factor of 14 from the slightly less-accurate [15].
-Acknowledgements.
+Time to solution for the TS is 1 500 000 core-hours (17 times faster than the benchmark [10] which had a simpler, grouped atom Jastrow but FCI input).
+6 Appendix: Jastrow factor and its structure.
+Jastrow factor: structure:
+The range is manifest in the value of ’cuttoff’ limited by half te lattice parameter of the supercell in real space.
+TERM1: e-e term (typically short range)
+The cutoff is energy is sensitive to its values. Two are required, for the range of electron 1 and electron 2. They are otherwise independent.
+In the TS runs for this work:
+Channel 1-1: L: [ 7.43797980979164, optimizable, limits:
+[ 0.500000000000, 10.4548873051022 ] ]
+Channel 1-2: L: [ 8.80054878236722, optimizable, limits:
+[ 0.500000000000000, 10.4548873051022 ] ]
+The polynomial expansion (in terms of r12) is of order 9.
+TERM 2:the e-n term (long range)
+Channel 1-n1: L: [ 9.00106155763939, optimizable, limits:
+[ 0.500000000000000,10.4548873051022 ] ]
+Channel 1-n2: L: [ 9.61549194785756, optimizable, limits:
+[ 0.500000000000000, 10.4548873051022 ] ]
+The polynomial expansion (in terms of r) is of order 9.
+Double sum, three particle (He-atom like) term. Over order 2 (e.g.4):
+TERM 3: Rank: [ 2, 1 ] e-e basis: [ Type: natural power, Order: 4 ] e-n basis: [ Type: natural power, Order: 4 ] e-n cutoff: Type: alt polynomial Constants: [ C: 3 ]
+Each atom has a specific range (usually long-range). Lowest in this case is just below 9 au. Mostly around 10. The value of a, lattice parameter is 21 au.
+Rules (restrictions on optimisation):
+Rules: [ 1-1=2-2, N, 1-N1=2-N1, 1-N2=2-N2, 1-N3=2-N3, 1-N4=2-N4, 1-N5=2-N5, 1-N6=2-N6, 1-N7=2-N7, 1-N8=2-N8, 1-N9=2-N9, 1-N10=2-N10, 1-N11=2-N11, 1-N12=2-N12, 1-N13=2-N13, 1-N14=2-N14, 1-N15=2-N15, 1-N16=2-N16, 1-N17=2-N17, 1-N18=2-N18, 1-N19=2-N19, 1-N20=2-N20, 1-N21=2-N21 ] QMC data on Figshare QMC Hoggan (Raw QMC data for AQC vol 92) Acknowledgements.
 We thank GENCI for a special allocation on Irene (nov 2022 to nov 2023 CEA, Bruyères-le-Châtel, F) and renewals until 2025, as well as University of Clermont Mesocentre (data preparation, tests).